# Shahriar Noroozizadeh

### v1

@@ 5 Related work
Our work consolidates fragments of a nascent phenomenon and weaves together distinct lines of theoretical and empirical work on memorization, learning compositional functions, and interpreting model representations. We elaborate on each of these threads below.
@@ 5.1 In-weights reasoning tasks
Synthetic graph tasks. Our work consolidates positive results of in-weights reasoning in literature, and presents a stronger instance of it, removing various confounders that make composition-learning easy. Khona et al. [76] report successful path-finding on $200$ nodes-large in-weights graphs, with varying path lengths and test-train overlap. Ye et al. [172], Wang et al. [158] report positive results on much shorter 2-hop tasks over $1000$ entities. Geerts et al. [43] look at in-weights transitive inference, a special type of $\ell$-fold composition query where the model is trained on local comparisons and is queried on more distant comparisons. On settings with $7$ objects, Geerts et al. [43] find a clear difference between an in-context version of this task (where the model struggles) and an in-weights version (where the model succeeds). Such relational queries are equivalent to giving two nodes along a path and querying which node is closer to the center of the graph. Our task of finding the first node is a much harder search task, where one finds the smallest node in an ordered relationship. Nagarajan et al. [108] discuss the limitations of next-token prediction, including on open-ended in-weights tasks, in lower data regimes. The fact that their next-token predictor achieves non-trivial performance on their in-weights task could be attributed to the effects of a geometric memory. Tangentially related is the positive finding in Yin and Wang [173] that the Transformer can compose in-weights knowledge given in-context demonstrations. It is worth noting that the (theoretical) arguments in both these works [108, 173] rest on the associative memory view.
@@
Multi-hop question-answering. A line of work has looked at natural language based two-hop questions on pretrained models e.g., “What is the calling code of the birthplace of Frida Kahlo?” (example from Press et al. [122]). Results here have been limited [122, 18] or mixed [169, 170, 14]. Yao et al. [171], Wang et al. [158] study multi-hop queries on synthetic knowledge (with the former on 2-hop queries while the latter explore upto 4 hops) and find that these queries can be learned provided there is exponential amounts of data, or a curriculum, or very long amounts of training. Balesni et al. [14] report that models are unable to compose synthetic facts, but can succeed in composing a synthetic fact with a natural one. Orthogonally, Wang et al. [162] identify that such in-weights implicit reasoning can be hurt by scaling up the parameters. Perhaps some of these negative results may be attributed to the reversal curse [15, 9], or the lack of extended computation e.g., we use pause tokens [52]. A dedicated study of this gap between our synthetic settings and these settings is important and left for future work.
Reversal curse and (a)symmetric knowledge. The reversal curse [15, 9] is a well-known out-of-distribution, in-weights failure mode of next-token-trained Transformers. Such models are unable to recall $u$ given $v$, when trained to recall $v$ given $u$, suggesting asymmetric storage in parametric memory. Fixes for the reversal curse have involved reversed or permuted data augmentation [57, 91, 79, 50]. In our settings, we find that reverse edges are critical to elicit implicit reasoning in our path-star tasks (see §D.1), but is not necessary to elicit a geometry in the smaller graphs. [^5] It appears that the observations on small graphs in Khona et al. [76] do not require memorizing the reverse edges, which aligns with our findings on small graphs. Various theories have been proposed to understand this failure [87, 183, 157], which may be worth revisiting under a geometric view. One may also view our contrast between associative and geometric memory (of a generic graph data) as a generalization of the aforementioned contrast between the asymmetric and symmetric knowledge storage (of a more specific, disjoint set of associations). We leave it for future work to discover a nuanced connection between our work and the reversal curse.
@@ 5.2 Failure of end-to-end composition learning
While $\ell$-fold composition functions are surprisingly easy to express in transformers [136], empirical results have time and again demonstrated that they are hard to learn through gradient-based methods, both in traditional deep network settings [141, 2, 55, 49, 1, 2] and more recently in language models too [112, 88, 30, 121, 176, 130, 30, 64, 145]. Others [12, 65, 142] demonstrate how next-token learning can trap training at a stage where composition learning becomes a problem. Theoretical works have attempted to formalize these failures by demonstrating limits due to to expressivity [97, 28, 119] or sample complexity [141] or computational complexity [165, 65] or in terms of statistical queries [164]. These results do not prove that a fixed composition function cannot be learned—only that a worst-case function exists for the given learning algorithm. Proving hardness for a fixed, singleton function class requires proving hardness of the full parity, a result that has only been recently proven [6, 144]. Finally, we note that all these results are concerned with composing in-context information; extending the negative results to composing in-weights information (within the associative view) likely requires non-trivial extensions, which we point out as an open theoretical question.
@@ 5.3 In-context graph tasks
While Bachmann and Nagarajan [12] identify the path-star topology as a failure case for next-token learning, Frydenlund [41, 42] demarcate the extent of this failure in the same in-context setting, whereas Brinkmann et al. [20] report positive path-finding results in other graph topologies. Others [137, 135] study other in-context graph search and counting tasks. Connections between in-context graph tasks and spectral biases exist [31, 116] but should not be confused with the spectral bias in in-weights tasks. While all these works study symbolic graph tasks, other works have empirically identified the limitations on graphs described in natural language [56, 159, 33]. Kim et al. [77], Ying et al. [175] propose algorithmic ideas for encoding graphs as inputs to Transformers. Finally, various failures [104, 35, 152, 153, 154, 143] have been reported on in-context tasks, including planning tasks, framed as word problems.
5.4 Analysis of Transformer memory
Associative memory. The concept of associative memory dates back to theories of how information is stored in the brain [90, 166]. These ideas have since been explicitly modeled through various architectures such as Hopfield networks [63, 129], energy-based models [80], and other modern Transformer-style inventions [82, 62]. Closest to us are work that analyze how architectures implicitly behave as associative memory storage, such as in autoencoders [126] or Transformers [17, 22, 111, 44, 139, 148]. We emphasize that this view is sufficient to understand Transformer behavior on disjoint facts, evidenced by the rich empirical literature built on this view. The geometric view only seems necessary when the facts become interdependent.
Empirical analyses. Other works [10, 105, 131, 115] have performed careful empirical analyses of scaling laws for memorization, and quantified memorization in terms of “bits per parameter count”, known as bit complexity [155], which is related to our notion of bit count in Proposition˜1. Zucchet et al. [184] empirically analyze the dynamics behind how facts are memorized in a model. Others [89, 178] have proposed methodological improvements to acquiring knowledge in a Transformer; of relevance to us is the finding in Zhang et al. [178] that training a model simultaneously on both facts and question-answering is a better way to integrate knowledge into the parameters.
@@ 5.4 Analysis of Transformer memory
5.5 Other foundational works on generalization memorization
Spectral and simplicity bias. The type of spectral bias we study in memorization and sequence modeling must be distinguished from the one studied in generalization and traditional classification and regression settings [127, 168, 74, 11, 133, 16]. In these earlier studies, the spectrum is that of a continuous function or a decision boundary (e.g., say, the Fourier components of a polynomial), whereas the spectrum we are concerned with is of a discrete, combinatorial object, namely the graph adjacency matrix. Furthermore, the core idea of these earlier studies is that the topmost eigenvectors (the lowermost frequencies) are learned first and the rest picked up later, suggesting the need to early-stop to preserve the top eigenvectors; whereas in our setting, longer training is required to filter out the bottom eigenvectors.
@@ 5.5 Other foundational works on generalization memorization
5.6 Analyses of graph and word embedding methods
Much attention has been given to characterizing what embeddings are learned by various contrastive losses such as Node2Vec. Most of these are on losses simpler than the softmax loss. However, a recent line of work on next-token prediction with the softmax loss [180, 179, 151] studies the geometry of Word2Vec models. The approach here is orthogonal to ours as they make connections to a support vector machine rather than the graph spectrum. The setting also has certain technical differences under which the training dynamics turn out to be very different e.g., the model converges in direction. The difference here likely stems from the fact that in these studies the embedding and unembedding matrices are not weight-tied and correspond to different spaces (words vs. contexts).
The connection to a graph spectrum has been made in many other analyses. These analyses focus on simpler loss called the negative sampling loss where the closed form expression for the inner products is straightforward (namely, the so-called pointwise mutual information (PMI) matrix, as discovered in Levy and Goldberg [84]). This analysis does not however tell us what exactly the embeddings are. The connection between these embeddings and the graph spectrum has been established in adjacent settings, like with DeepWalk [124] (where the objective is explicitly multi-hop), in low-rank settings like SimCLR [59, 149] or with quadratic losses with early stopping [75] or the softmax loss with rank $1$ [68]. Other analyses of Node2Vec focus on specific types of graphs such as stochastic block models [34, 60]. We refer the reader to Goyal and Ferrara [51] for a survey of graph embedding methods. Finally, we clarify that these methods and our insights must not be confused with graph neural networks [174, 181], where the graphs are not stored in parametric memory, but presented as input.
@@ 6 Limitations
    We emphasize that all our arguments (e.g., about the lack of pressure) are empirical and informal. Perhaps, a slightly more nuanced form of architectural or statistical pressure (e.g., more nuanced norm complexity such as flatness of the loss) may indeed explain why associative memory is less preferred by the model. Conversely, perhaps the lack of pressures in the learning setup is indeed why the Transformer learns a sub-optimal kind of geometric memory compared to Node2Vec.
@@ 7 Conclusion
Various practical questions arise out of these findings. The elegant geometries of Node2Vec models indicate that Transformer memory can be made much more geometric. It is also unclear whether Transformers exhibit such desirable behaviors in more complex graph topologies. Importantly, empirical works on implicit reasoning in natural language have so far been mixed [122, 18, 169, 170, 14]. More careful empirical research and ideation may be needed to make the geometric view more broadly applicable. Orthogonally, our findings are of relevance to making choices between parametric vs. contextual memory, and also between generative retrieval vs. dual encoder retrieval models.
Our work raises the foundational question of when and how associative and geometric memory compete with each other during optimization, and what factors—such as training time, learning rate, weight decay—can foster one over the other. To answer these, our insights into the spectral bias need to be extended from Node2Vec-style architectures (where associative memory is prohibited) to deep sequence models (where associative memory becomes a competitor). We also hope that our simple examples can be useful for orthogonal conceptual studies on interpretability, world models, and in characterizing convergent representations across model families. Broadly, we expect these findings to inspire revisiting unstated associative assumptions underlying research on knowledge and memory in language models.
Acknowledgments. We would like to thank Gaurav Ghosal, Gintare Karolina Dziugaite, George H Chen, Christina Baek, and Jacob Springer for valuable feedback on earlier versions of this draft. We are also grateful to Andrej Risteski for discussions.

### v2

@@ 5 Related work
Our work consolidates fragments of a nascent phenomenon and brings together distinct lines of theoretical and empirical work on memorization, learning compositional functions, and interpreting model representations. We elaborate on each of these threads below.
5.1 Analysis of Transformer memory
Associative memory. The concept of associative memory dates back to theories of how information is stored in the brain [101, 182]. These ideas have since been explicitly modeled through various architectures such as Hopfield networks [70, 142], energy-based models [89], and other modern Transformer-style inventions [91, 69]. Closest to us are works that analyze how architectures implicitly behave as associative memory storage, such as in autoencoders [139] or Transformers [20, 26, 124, 49, 154, 164]. Others have established connections between the attention mechanism and Hopfield networks [142, 161]. We emphasize that this view is sufficient to understand Transformer behavior on disjoint facts, evidenced by the rich empirical literature built on this view. The geometric view only seems necessary when the facts become interdependent.
Geometric memory as node representations. Although, geometric memory has not been explicitly labeled as a distinct form of parametric memory, we find scattered observations of it in literature in weaker forms. [78, 85, 189] show that nodes in graph-like datasets are embedded in a geometric way that encodes distances. These settings involve some form of global supervision or latent high-level structure, while we find that such a geometry can arise even with atomic facts. Jiang et al. [78] further juxtapose these structured embeddings against associative memory. However, they suggest that the structure becomes relevant only in the underparameterized setting, and that an associative memory suffices as an abstraction in the overparameterized setting; we find this is not the case. Independently, on two-hop datasets, [189, 174] report how models first “memorize” the training data, before “generalizing” on their reasoning task; under the hood, this translates to associative and geometric memorization respectively.
Geometric memory as factorized knowledge. Recall that the geometric memory can be viewed as storing a low-rank factorization of the adjacency matrix. The advantages of such factorizations is again scattered in literature. Saxe et al. [153, 152] prove that a two-layer model naturally learns an SVD factorization of input-output associations. This is a powerful result demonstrating the value of depth in learning hierarchical representations; but the global, multi-hop nature of these representations was not highlighted. Also, unlike our setting, theirs requires explicit pressures to demonstrate a low rank spectral bias (either through a bottleneck or early-stopping) for they study the squared-error loss. [^10] Under their squared error loss (applied without a softmax layer), a fully-trained 2-layered model would learn the 1-hop adjacency matrix perfectly, with no global, multi-hop information revealed in the logits. Although the associations are stored in a factorized form, the final hidden representations do not filter out the lower directions. More recently, Huang et al. [74] point out that two-hop reasoning emerges if and only if the key-query matrices in an attention layer are kept separate (as against being collapsed into a single matrix); we show that such favorable effects occur in more general architectures without attention layers. They also highlight the role of an implicit regularization effect from their reasoning supervision; we emphasize that this does not resolve our memorization puzzle—see Remark˜1.
We also interpret these prior analyses [153, 74] as a study of how $2$-layered models memorize a special type of graphs: bipartite graphs (although they do not express it as such). [^11] Concretely, [153] study mappings between “entity” tokens and binary feature vectors; we translate this as a bipartite graph where “entity” nodes are connected to various “feature” nodes. [74] study a mapping between “subject” nodes and “answer” nodes. Importantly, as noted in our own analysis in §4, such 2-layered models do not capture the competition between associative and geometric memory. [^12] At least not in its full glory; in $2$-layered weight-tied models, there is only a contrived form of associative memory (Proposition 5); if weight-untied, a natural form can arise in a single step (Proposition 4). For that, a study of deeper models is required, where the dynamics along the various spectral components become entangled; to keep them untangled, unnatural assumptions about the initializations need to be made, as shown in [152]. Another nuance in these prior 2-layered analyses is the lack of weight-tying, which creates significant differences in what is learned, as we identify in §4.4.
Remark 1.
(Implicit regularization in matrix factorization is an insufficient explanation of geometric memory) Huang et al. [74] attribute the success of reasoning to implicit regularization effects that arise in matrix completion tasks; we argue that these effects do not explain why a geometry arises over associative storage in our experiments. The “matrix completion” view in [74] is that during training the model attempts to fit certain observed entries in a $2$-hop adjacency matrix ${\bf{A}}^{2}$; any test query in effect demands that the model “complete” a missing $2$-hop entry. While many minima exist in such underdetermined matrix completion tasks, it is known [62, 11] that gradient descent on a deep factorized model finds the least nuclear norm matrix consistent with the observed entries. However, our edge-memorization-only task is a well-determined task that requires storing the adjacency matrix $\bf{A}$} as is, without any missing entries to infer; thus no such implicit pressure must exist. Yet a geometry arises in this task as shown in Section˜3.1. While it is still possible that there is a certain nuclear norm minimization effect even in edge-memorization tasks—perhaps because the logits themselves are not well-determined—the connection is not as obvious as it may seem.
Empirical analyses. Other works [10, 115, 144, 128] have performed careful empirical analyses of scaling laws for memorization, and quantified memorization in terms of “bits per parameter count”, known as bit complexity [171], which is related to our notion of bit count in Section˜3.2. All these analyses are worth revisiting with the newfound lens of a geometric memory. Zucchet et al. [203] empirically analyze the dynamics behind how facts are memorized in a model. Others [99, 196] have proposed methodological improvements to acquiring knowledge in a Transformer; of relevance to us is the finding in Zhang et al. [196] that training a model simultaneously on both facts and question-answering is a better way to integrate knowledge into the parameters.
@@ 5.2 In-weights reasoning tasks
Synthetic graph tasks. Our work consolidates positive results of in-weights reasoning in literature, and presents a stronger instance of it, removing various confounders that make composition-learning easy. Khona et al. [85] report successful path-finding on $200$ nodes-large in-weights graphs, with varying path lengths and test-train overlap. Others [174, 43, 189, 74] report positive results on much shorter 2-hop tasks over a thousand entities or fewer. Geerts et al. [48] look at in-weights transitive inference, a special type of $\ell$-fold composition query where the model is trained on local comparisons and is queried on more distant comparisons. On settings with $7$ objects, Geerts et al. [48] find a clear difference between an in-context version of their task (where the model struggles) and an in-weights version (where the model succeeds). Our task of finding the first node is a much harder search task akin to finding the smallest node in an ordered relationship. Nagarajan et al. [119] discuss the limitations of next-token prediction, including on open-ended in-weights tasks, in lower data regimes. The fact that their next-token predictor achieves non-trivial performance on their in-weights task could be attributed to the effects of a geometric memory. Tangentially related is the positive finding in Yin and Wang [191] that the Transformer can compose in-weights knowledge given in-context demonstrations. It is worth noting that the (theoretical) arguments in both these works [119, 191] rest on the associative memory view.
@@
Multi-hop question-answering. A line of work has looked at how pretrained models respond to two-hop questions in natural language e.g., “What is the calling code of the birthplace of Frida Kahlo?” (example from Press et al. [135]). Results here have been limited [135, 21] or mixed [185, 186, 16]. Success has been shown on synthetic knowledge provided there is exponential amounts of data, or a curriculum, or very long amounts of training [188, 174]. Orthogonally, Wang et al. [178] identify that such in-weights implicit reasoning can be hurt by scaling up the parameters, perhaps due to “excessive memorization”, which we interpret as associative memorization. It is possible that some of these negative results may be attributed to the reversal curse [18, 9], or the lack of extended computation e.g., we use pause tokens [57]. But a dedicated study of this gap between our synthetic settings and these settings is important and left for future work.
Reversal curse and (a)symmetric knowledge. The reversal curse [18, 9] is a well-known out-of-distribution, in-weights failure mode of next-token-trained Transformers. Such models are unable to recall $u$ given $v$, when trained to recall $v$ given $u$, suggesting asymmetric storage in parametric memory. Fixes for the reversal curse have involved reversed or permuted data augmentation [64, 102, 88, 55]. In our settings, we find that reverse edges are critical to elicit implicit reasoning in our large path-star tasks (see §D.1), but is not necessary to elicit a geometry in the smaller graphs. [^13] It appears that the observations on small graphs in Khona et al. [85] do not require memorizing the reverse edges, which aligns with our findings on small graphs. Various theories have been proposed to understand this failure [97, 202, 173], which may be worth revisiting under a geometric view. One may also view our contrast between associative and geometric memory (of a generic graph data) as a generalization of the aforementioned contrast between the asymmetric and symmetric knowledge storage (of a more specific, disjoint set of associations). Future work may discover a nuanced connection between our work and the reversal curse.
@@ 5.3 Failure of end-to-end composition learning
While $\ell$-fold composition functions are surprisingly easy to express in transformers [149], empirical results have time and again demonstrated that they are hard to learn through gradient-based methods, both in traditional deep network settings [156, 2, 61, 54, 1, 2] and more recently in language models too [125, 98, 34, 134, 194, 143, 34, 71, 160]. Others [13, 72, 157] demonstrate how next-token learning can trap training at a stage where composition learning becomes a problem. Theoretical works have attempted to formalize these failures by demonstrating limits due to expressivity [108, 32, 132] or sample complexity [156] or computational complexity [181, 72] or in terms of statistical queries [180].
Remark 2.
(Theoretical hardness of learning compositions) Existing theoretical hardness results typically only show that some worst-case function in a class of compositional functions is hard-to-learn; no proclamations are made about how hard it is to learn a fixed function we care about. Indeed, with contrived initial conditions, a singleton function class can be provably learned [4, 3, 118]. Yet, in practice, learning a fixed function is hard, proving which has been an open question—until the recent negative results of Abbe et al. [6], Shoshani and Shamir [159]. These show how the (fixed) full parity function is hard to learn with gradient descent which, through a reduction, proves that in-context composition tasks are hard to learn [72]. We leave it as an open question to link our in-weights composition task to one of these hardness results.
@@ 5.4 In-context graph tasks
While Bachmann and Nagarajan [13] identify the path-star topology as a failure case for next-token learning, Frydenlund [46, 47] demarcate the extent of this failure in the same in-context setting, whereas Brinkmann et al. [24] report positive path-finding results in other graph topologies. Others [150, 148, 190] study Transformers in various in-context graph tasks like search and counting. Connections between in-context graph tasks and spectral biases exist [35, 129] but should not be confused with the spectral bias in in-weights tasks. While all these works study symbolic graph tasks, other works have empirically identified the limitations on graphs described in natural language [63, 175, 37]. Kim et al. [86], Ying et al. [193] propose algorithmic ideas for encoding graphs as inputs to Transformers. Finally, various failures [114, 39, 168, 169, 170, 158] have been reported on in-context tasks, including planning tasks, framed as word problems.
@@
5.5 Analyses of graph and word embedding methods
Much attention has been given to characterizing what embeddings are learned by various contrastive losses such as Node2Vec. Most of these are on losses simpler than the cross-entropy loss, with the exception of a recent line of work [199, 198, 167] which studies the geometry of cross-entropy-trained (weight-untied) Word2Vec models. However, certain assumptions are made, such as freezing one of the layers [167] or heuristically approximating via a regularized proxy model [199]. Curiously, [199] find that their system only directionally convergences, while the weights themselves diverge; we however point out that it’s reasonable to expect the model to converge to a zero-gradient solution. Regardless, [199] suggest an insightful connection to the converged direction to support vector machines; this may inspire better characterizations of what solution is found by our unregularized Node2Vec system.
The connection to a graph spectrum has been made in many other analyses. These analyses focus on a simpler loss called the negative sampling loss where the closed form expression for the inner products is straightforward (namely, the so-called pointwise mutual information (PMI) matrix, as discovered in Levy and Goldberg [93]). This analysis does not however tell us what exactly the embeddings are. The connection between these embeddings and the graph spectrum has been established in adjacent settings, like with DeepWalk [137] (where the objective is explicitly multi-hop), in low-rank settings like SimCLR [66, 165] or with quadratic losses with early stopping [83] or the softmax loss with rank $1$ [76]. Other analyses of Node2Vec focus on specific types of graphs such as stochastic block models [38, 67]. We refer the reader to Goyal and Ferrara [56] for a survey of graph embedding methods. Finally, we clarify that these methods and our insights must not be confused with graph neural networks [192, 200], where the graphs are not stored in parametric memory, but presented as input.
5.6 Other foundational works on generalization and memorization
Spectral and simplicity bias. The type of spectral bias we study in memorization and sequence modeling must be distinguished from the one studied in generalization and traditional classification and regression settings [140, 184, 82, 12, 146, 19]. In these earlier studies, the spectrum is that of a continuous function or a decision boundary (e.g., say, the Fourier components of a polynomial), whereas the spectrum we are concerned with is of a discrete, combinatorial object, namely the graph adjacency matrix. This corresponds to a difference in the nature of representations as remarked below.
Remark 3.
(On the nature of representations in classification vs. sequence modeling) The geometric representations learned in our setting (i.e., sequence modeling of atomic facts) are of a fundamentally different nature from that learned in the classification/regression setting. The representations in classification/regression correspond to features inherent to each datapoint, such as the edges detected in an image. No such features are inherent to a one-hot token; a token’s representation is synthesized from its co-occurrence patterns with other tokens. This difference is central to why the analogy between the generalization puzzle and the memorization puzzle breaks down in §3.2, both in terms of succinctness and ease of discovery with gradient descent.
@@ 6 Limitations
    We emphasize that all our arguments are empirical and informal (or left as propositions), and lack a precise, fuller characterization of our observations.
    1. (a)
      Perhaps, a slightly more nuanced form of architectural or statistical pressure (e.g., more nuanced norm complexity such as flatness of the loss) may indeed explain why associative memory is less preferred by the model.
    2. (b)
      Conversely, perhaps the lack of pressures in the learning setup is indeed why the Transformer learns a sub-optimal kind of geometric memory compared to Node2Vec.
@@ 7 Conclusions and Future Work
Various practical questions arise out of these findings. The elegant geometries of Node2Vec models indicate that Transformer memory can be made much more geometric. It is also unclear whether Transformers exhibit such desirable behaviors in more complex graph topologies. Importantly, empirical works on implicit reasoning in natural language have so far been mixed. More careful empirical research and ideation may be needed to make the geometric view more broadly applicable. Orthogonally, our findings are of relevance to making choices between parametric vs. contextual memory, and also between generative retrieval vs. dual encoder retrieval models.
Our work raises the foundational question of when and how associative and geometric memory compete with each other during optimization, and what factors—such as graph connectivity, training time, learning rate, weight decay—can foster one over the other. To answer these, our insights into the spectral bias need to be extended from Node2Vec-style architectures (where associative memory is prohibited) to deep sequence models (where associative memory becomes a competitor). We also hope that our simple examples can be useful for orthogonal conceptual studies on interpretability, world models, and in characterizing convergent representations across model families. Broadly, we expect these findings to inspire revisiting unstated associative assumptions underlying research on knowledge and memory in language models.
Acknowledgments. We would like to thank Gaurav Ghosal, Gintare Karolina Dziugaite, George H Chen, Christina Baek, Jacob Springer and Omar Salemohamed for valuable feedback on earlier versions of this draft. We are also grateful to Andrej Risteski, Alberto Bietti, Ekdeep Singh Lubana and Andrew Lampinen for discussions and pointers to related work.